OpenStudy (anonymous):

line passing through (9,2) and (-2,6)

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

use y_2-y_1 / x_2-x_1

OpenStudy (lgbasallote):

\[\large y - y_1 = \frac{y_2 - y_1}{x_2 - x_1} (x - x_1)\]

OpenStudy (lgbasallote):

\[\large y - 2 = \frac{6-2}{-2 - 9} (x - 9)\] just simplify it